OpenStudy (anonymous):

What is the y-intercept of the line with the equation 4x + 9y = -108

OpenStudy (anonymous):

To find y-intercept, let x=0 and solve for y

OpenStudy (gw2011):

Another way of looking at this is to use the equation of a line which is: y = mx +b where m is the slope and b is the y-intercept Therefore, 4x + 9y = -108 9y = -4x - 108 y = -(4/9)x - 108/9 y = -(4/9)x - 12 Therefore, the y-intercept is -12